The balance sheet for Shrubs Landscaping, Inc. includes

Cash $54,000

Short-term investments $154,000

Accounts Receivable $96,000

Inventory $360,000

Prepaid Insurance $42,000

Equipment $675,000

Accounts Payable $150,000

Unearned Revenue $100,000

Salaries Payable $15,000

Taxes Payable $5,000

Bonds Payable $750,000

Stockholder’s Equity $361,000

Calculate the acid-test (quick) ratio for Shrubs: (Round your answer to two decimal places.)

2.46

2.61

2.03

1.13
Answer #1

Acid test ratio = Quick assets/Current liabilities

Quick assets = 54,000+154,000+96,000 = 304,000

Current Liabilities = 150,000+100,000+15,000+5,000 = 270,000

= 304,000/270,000

= 1.13

Option D is the answer